OpenStudy (anonymous):

solve for X: 2/x+2 + 1/5 = 6/x+5

OpenStudy (anonymous):

Thats coool

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

what school ?

OpenStudy (anonymous):

A school In Queens..Why?

OpenStudy (anonymous):

Just wondering, never met someone with the last name as me

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

-10/7

OpenStudy (anonymous):

if i understand it correctly it's (2/x)+2 + 1/5 = (6/x)+5 , then x = - 10/ 7

OpenStudy (anonymous):